In a discount inter- est loan, you pay the interest payment up front. For example, if a 1-year loan is stated as $10,000 and the interest rate is 10%, the borrower “pays” .10 × $10,000 = $1,000 immediately, thereby receiving net funds of $9,000 and repaying $10,000 in a year.

I have this question and i know the answer is 11.11 but im not sure how you came to this answer.

i got 10,000=9,000 x (1+r) but im not getting the right answer when I do this.

help please
Answer #1

You got the equation right, you may have made error in steps while solving it:

Answer:

Let us assume the effective interest rate is = r

Hence from information given, we have:

10000 = 9000 * (1+r)

=> 9000 * (1+r) = 10000

=> 9000 + 9000* r = 10000

=> 9000* r = 10000 - 9000

=> 9000* r = 1000

=> r = 1000 / 9000

=> r = 0.1111 or 11.11%

So the answer is 11.11%